# Vision-driven River Following of UAV via Safe Reinforcement Learning using Semantic Dynamics Model

### v1

@@ 2.4 Model-based Safe Reinforcement Learning
In safe reinforcement learning (SafeRL), the objective is to optimize a policy while keeping the cumulative cost below a certain threshold [7]. Decoupling task-specific rewards from environment-driven costs ensures that safety remains an independent priority rather than being tied to task performance. Many model-free RL algorithms [18, 19, 36, 37, 38, 39, 40] try to solve the safe optimal policy in CMDP using the Lagrangian method. However, in model-free Lagrangian SafeRL, suboptimal performance under low-cost thresholds arises because the critic’s estimation inaccuracies lead to frequent constraint violations, as the agent lacks precise cost information for its actions [17]. Besides, safe model-free algorithms still require extensive interactions with the environment to converge, while model-based methods can significantly enhance sample efficiency by also learning from the imagined trajectories [37]. Moreover, as pointed out in [41, 42], system dynamics knowledge, either learned from data or acquired from prior modeling, is necessary for stronger safety guarantees.
@@
Building on the better data efficiency and constraint satisfaction of model-based approaches, several safety regulation paradigms have incorporated dynamics models to enhance SafeRL. First, Control Barrier Functions (CBFs) enforce one-step safety by solving a quadratic program that constrains the Lie derivative of a smooth barrier function, thereby guaranteeing forward invariance of a safe set [43, 44, 45, 46, 47, 48]. However, synthesizing this function requires domain knowledge and does not scale to high-dimensional visual observations [49]. Shielding methods extended this idea by pre-computing backup policies or reachable safe sets that logically block unsafe actions at runtime [50, 51, 52, 53], yet they depend on formal specifications and generally do not feed back into policy learning. Third, model-based Lagrangian policy optimization techniques [37, 17, 36, 54] embed cost constraints directly into the policy gradient objective via dual variables, allowing the agent to learn trade-offs between cumulative reward and cost. However, Lagrangian-based methods ensure constraint satisfaction in expectation during training, but do not enforce per-step safety once the policy is deployed. Fourth, proactive model predictive shielding (MPS) approaches, including Statistical MPS [55] and Dynamic MPS [56, 57], either uses probabilistic check to ensure safety with high probability, or rank candidate actions using finite-horizon rollouts through a known or learned dynamics model to dynamically synthesize shielding action at each step, co-training the planner with the policy to improve both safety and performance.
@@
In comparison to classic CBF or shielding methods that project or reject unsafe actions as post-hoc safety filters without altering the underlying policy, we focus on using a model predictive safety shield that regulates policy during training like Lagragian-based methods, but also can be activated during inference like MPS approaches. While existing Lagrangian-based or MPS methods incorporate dynamics models for safety-aware planning, they primarily rely on cost critics to estimate long-term safety violations. This is not the case in our UAV river/track waypoint navigation setting, where the vehicle is treated as point mass that can simply reverse course upon sensing danger. Thus, the genuine safety failures only occur the moment of, for example collision, not from drifting too far into the future [58, 59]. Consequently, immediate Markovian costs, captured by a cost estimator, are preferable to provide timely danger signals for policy updates, without the complexity or potential inaccuracy of long-horizon cost value functions, for agents with ego-centric observations. However, leveraging predicted immediate costs to form a cost advantage for SafeRL, or incorporating this signal into a safety layer during training, remains underexplored.
@@ 3.1 Constrained Submodular Markov Decision Process
$\pi^{\ast}=\arg\max_{\pi\in\Pi_{C}}J^{R}(\pi)$ (1)
A policy’s optimality is determined by its performance measure $J^{R}(\pi)\in\mathbb{R}_{\geq 0}$, which is usually the expected discounted cumulative rewards over the infinite horizon: $J^{R}(\pi)\doteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{\pi}}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$. A policy’s feasibility is determined by its safety measure $J^{C}(\pi)\in\mathbb{R}_{\geq 0}$, which is similar as the performance measure in definition: $J^{C}(\pi)\doteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{\pi}}[\sum_{t=0}^{\infty}\gamma^{t}C(s_{t},a_{t})]$. Thus $\Pi_{C}\doteq\{\pi\mid J^{C}(\pi)\leq d\}$ is the feasible policy set, where $d$ is a cost budget hyperparameter. In this work we only focus on single cost constraint, but CMDP does not have restrictions on the number of cost constraints. In CMDP, the reward and cost advantage functions are both defined as the difference between the state action value function $Q_{\pi}(s,a)\doteq E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a]$ and the state value function $V_{\pi}(s)\doteq E_{\tau\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s]$, i.e., $A_{\pi}(s,a)=Q_{\pi}(s,a)-V_{\pi}(s)$. However, in environments where the reward function is submodular and non-Markovian, the traditional state-wise advantage estimation may fail to capture the trajectory-dependent nature of rewards, leading to suboptimal policy updates. Therefore, a trajectory-wise advantage estimation that considers the cumulative submodular rewards is necessary for more effective learning in these complex environments.
A Constrained Submodular Markov Decision Process (CSMDP) extends the CMDP by recognizing the submodularity of the reward function, meaning the reward in CSMDP is non-additive and non-Markovian (history-dependent), and its value decreases in light of similar states visited previously, [6]. Formally, the set function $R$ in CSMDP can be denoted as $R_{sub}$, which is submodular. For two subsets $S_{1},S_{2}$ of set $S$ that meet $\forall S_{1}\subseteq S_{2}\subseteq S,s\in S\setminus B$, we have
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{sub}(S_{1}\cup s)-R_{sub}(S_{1})\geq R_{sub}(S_{2}\cup s)-R_{sub}(S_{2})$} (2)
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta(s|S_{1}):=R_{sub}(S_{1}\cup s)-R_{sub}(S_{1})$} (3)
$R_{sub}$ is a monotone submodular function if $R_{sub}(S_{1})\leq R_{sub}(S_{2}),\forall S_{1}\subseteq S_{2}\subseteq S$. In plain words, in CSMDP, adding an element $s$ to the subset $A$ will help at least as much as adding it to the superset $B$, in terms of rewards gaining. Besides, the marginal gain $\Delta(s|A)$ will be non-negative when the reward function $R_{sub}$ is monotone submodular. This is exactly the case in track/river following tasks in environments like CliffCircular [4] and Safe Riverine Environment [20], where repeated visiting of the visited states will not bring neither positive nor negative rewards. The punishments to suboptimal policies will be reflected by the cost function $C(s,a)$. It is critical for safe RL agents in CSMDP to understand the submodular and non-Markovian nature of the reward function, leverage this knowledge, and optimize a policy that effectively balances reward maximization with cost minimization.
@@ 3.2 Marginal Gain Advantage Estimator
where $\bar{D}_{KL}(\pi||\pi_{k})=E_{s\sim\pi_{k}}[\bar{D}_{KL}(\pi||\pi_{k})[s]]$, $D_{KL}(\pi||\pi_{k})$ is the Kullback–Leibler divergence between the target unparameterized policy $\pi$ and the current parameterized policy $\pi_{k}$, $\delta$ is the target KL divergence, $\gamma$ is the discount factor, $i$ is the cost constraint index number ($i\in\{1\}$ in our experiments, single cost is considered).
@@
To project the optimal update policy $\pi^{*}$ to the parameterized policy $\pi_{\theta}$, let $\tau=(s_{0},a_{0},s_{1},a_{1},...,s_{T-1},a_{T-1})$ represent a trajectory of length $T$, then minimize the loss function trajectory-wise instead of state-wise,
$\mathcal{L}(\theta)=\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[D_{KL}\left(\pi_{\theta}\|\pi^{*}\right)[\tau]\right],$ (8)
@@
$\nabla_{\theta}\mathcal{L}(\theta)=\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[\nabla_{\theta}D_{KL}\left(\pi_{\theta}\|\pi^{*}\right)[\tau]\right]$
$=\nabla_{\theta}D_{KL}\left(\pi_{\theta}\|\pi_{\theta_{k}}\right)[\tau]$-
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{\alpha}}\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[\sum_{i=0}^{T-1}\frac{\nabla_{\theta}\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}\left(A_{R}^{\pi_{\theta_{k}}}(s_{i},a_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}A_{C}^{\pi_{\theta_{k}}}(s_{i},a_{i})\right)\right]$ (9)
We refer readers to [19] for detailed proof. The reward advantage $A_{R}^{\pi_{\theta_{k}}}(s_{i},a_{i})$ is defined in a submodularity-inspired trajectory-wise way, named Marginal Gain Advantage Estimation (MGAE) for reward,
@@
where $\Delta$ represents the true marginal gain of the state $s_{j+1}$ given the trajectory $\tau_{0:j}$ up to horizon $j$, and $b(\tau_{0:i})$ is the baseline function dependent on the trajectory up to horizon $i$ that helps to reduce the variance of reward advantage estimation [60]. The important part here is how to choose the baseline function. It should represent a reasonable anticipation of the future cumulative rewards that can guide the policy improvement by comparing with the actual partial return to form the reward advantage. As proved in [6], baseline can be any history-dependent function. We choose this baseline function to be the agent’s current best guess of the remaining rewards along the current trajectory to further reduce the variance,
$b(\tau_{0:i})=\underset{\tau\sim\pi_{\theta}}{\operatorname{mean}}({\sum_{i=0}^{T-1}\Delta(s_{i+1}\mid\tau_{0:i})})-\sum_{i=0}^{j}\hat{\Delta}(s_{i+1}\mid\tau_{0:i})$ (11)
The left part in Equation 11 is the mean trajectory return across a sliding window (window size is chosen as $10$ in both environments) of trajectories up to the baseline calculation ($\tau\sim\pi_{\theta}$), whereas the right part is the sum of estimated marginal gains up to step $j$ along the current trajectory. $\hat{\Delta}(s_{i+1}\mid\tau_{0:i})$ is the estimated marginal gain of the state $s_{i+1}$ given the trajectory $\tau_{0:i}$ up to horizon $i$. The illustrative comparison between MGAE and the popular GAE [8] is visually presented in Figure 1.
@@ 3.3 Semantic Dynamics Model
To further improve spatial consistency, our method employs a geometrically constrained semantic dynamics model (SDM), where a neural network estimates the homography between consecutive semantic observations. This approach assumes that a significant portion of pixels in the image represents a flat surface (e.g., water) in the 3D environment, allowing the homography to approximate the geometric transformation between frames effectively. Additionally, the model assumes that most objects in the scene are static, which helps simplify the estimation process. By preserving the spatial transformations under these assumptions, our method enables the RL model to better capture short-term dynamic changes in the environment. By integrating this structured approach, we provide a more semantically rich and geometrically coherent input to the RL model, which enhances not only the effectiveness of learning and planning but also contributes to safer decision-making by improving the predictability of future states and ensuring consistency in sequential observations.

### v2

@@ 2.4 Model-based Safe Reinforcement Learning
In safe reinforcement learning (SafeRL), the objective is to optimize a policy while keeping the cumulative cost below a certain threshold [8]. Decoupling task-specific rewards from environment-driven costs ensures that safety remains an independent priority rather than being tied to task performance. Many model-free RL algorithms [19, 20, 38, 39, 40, 41] try to solve the safe optimal policy in CMDP using the Lagrangian method. However, in model-free Lagrangian SafeRL, suboptimal performance under low-cost thresholds arises because the critic’s estimation inaccuracies lead to frequent constraint violations, as the agent lacks precise cost information for its actions [18]. In addition, safe model-free algorithms still require extensive interactions with the environment to converge, while model-based methods can significantly improve sample efficiency by also learning from imagined trajectories [42]. Moreover, as pointed out in [43, 44], system dynamics knowledge, either learned from data or acquired from prior modeling, is necessary for stronger safety guarantees.
@@
Building on the better data efficiency and constraint satisfaction of model-based approaches, several safety regulation paradigms have incorporated dynamics models to enhance SafeRL. First, Control Barrier Functions (CBF) enforce one-step safety by solving a quadratic program that constrains the Lie derivative of a smooth barrier function, thus guaranteeing forward invariance of a safe set [45, 46, 47, 48, 49, 50]. However, synthesizing this function requires domain knowledge and does not scale to high-dimensional visual observations [51]. Shielding methods extended this idea by precomputing backup policies or reachable safe sets that logically block unsafe actions at runtime [52, 53, 54, 55], yet they depend on formal specifications and generally do not feed back into policy learning. Third, model-based Lagrangian policy optimization techniques [42, 18, 38, 56] embed cost constraints directly into the policy gradient objective via dual variables, allowing the agent to learn trade-offs between cumulative reward and cost. However, Lagrangian-based methods ensure constraint satisfaction in expectation during training, but do not enforce per‐step safety once the policy is deployed. Fourth, proactive model predictive shielding (MPS) approaches, including Statistical MPS [57] and Dynamic MPS [58, 59], either uses probabilistic check to ensure safety with high probability, or rank candidate actions using finite-horizon rollouts through a known or learned dynamics model to dynamically synthesize shielding action at each step, co-training the planner with the policy to improve both safety and performance.
@@
In comparison to classic CBF or shielding methods that project or reject unsafe actions as post-hoc safety filters without altering the underlying policy, we focus on using a model predictive safety shield that regulates policy during training like Lagragian-based methods, but also can be activated during inference like MPS approaches. While existing Lagrangian-based or MPS methods incorporate dynamics models for safety-aware planning, they primarily rely on cost critics to estimate long-term safety violations. This is not the case in our UAV river/track waypoint navigation setting, where the vehicle is treated as point mass that can simply reverse course upon sensing danger. Thus, the genuine safety failures only occur the moment of, for example collision, not from drifting too far into the future [60, 61]. Consequently, immediate Markovian costs, captured by a cost estimator, are preferable to provide timely danger signals for policy updates, without the complexity or potential inaccuracy of long‐horizon cost value functions, for agents with ego-centric observations. However, leveraging predicted immediate costs to form a cost advantage for SafeRL, or incorporating this signal into a safety layer during training, remains underexplored.
@@ 3.1 Constrained Submodular Markov Decision Process
$\pi^{\ast}=\arg\max_{\pi\in\Pi_{C}}J^{R}(\pi).$ (1)
[…]
A Constrained Submodular Markov Decision Process (CSMDP) extends the CMDP by recognizing the submodularity of the reward function, meaning the reward in CSMDP is non-additive and non-Markovian (history-dependent), and its value decreases in light of similar states visited previously, [7]. Formally, the set function $R$ in CSMDP can be denoted as $R_{sub}$, which is submodular. For two subsets $S_{1},S_{2}$ of set $S$ that meet $\forall S_{1}\subseteq S_{2}\subseteq S,s\in S\setminus S_{2}$, we have
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{sub}(S_{1}\cup\{s\})-R_{sub}(S_{1})\geq R_{sub}(S_{2}\cup\{s\})-R_{sub}(S_{2}).$} (2)
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta(s|S_{1}):=R_{sub}(S_{1}\cup\{s\})-R_{sub}(S_{1}).$} (3)
Given a partial trajectory $\tau_{0:j}=(s_{0},a_{0},...,s_{j})$, let $S_{j}=\{s_{0},...,s_{j}\}$ denote the visited set. We write $\Delta(s_{j+1}\mid\tau_{0:j})$ as shorthand for $\Delta(s_{j+1}\mid S_{j})=R_{sub}(S_{j}\cup\{s_{j+1}\})-R_{sub}(S_{j})$. $R_{sub}$ is a monotone submodular function if $R_{sub}(S_{1})\leq R_{sub}(S_{2}),\forall S_{1}\subseteq S_{2}\subseteq S$. Here, $R_{sub}(S)$ denotes the submodular, set-based reward in the Submodular MDP, the cumulative reward accrued over the visited set $S$. The symbol $R(s,a)$ appears only in the preliminaries as standard MDP notation and is not used subsequently. In plain words, in CSMDP, adding an element $s$ to the sub-subset $S_{1}$ will help at least as much as adding it to the subset $S_{2}$, in terms of rewards gaining. Besides, the marginal gain $\Delta(s|S_{1})$ will be non-negative when the reward function $R_{sub}$ is monotone submodular. This is exactly the case in track/river following tasks in environments like CliffCircular [4] and Safe Riverine Environment [21], where repeated visiting of the visited states will not bring neither positive nor negative rewards. Safety is modeled separately via the cost signal $C(s,a)$. In the constrained formulation, the goal is to exploit the history-dependent marginal-gain structure of reward while ensuring the cost constraint is satisfied, Equation 1.
@@ 3.2 Marginal Gain Advantage Estimator
where $\bar{D}_{KL}(\pi||\pi_{k})=E_{s\sim\pi_{k}}[\bar{D}_{KL}(\pi||\pi_{k})[s]]$, $D_{KL}(\pi||\pi_{k})$ is the Kullback–Leibler divergence between the target unparameterized policy $\pi$ and the current parameterized policy $\pi_{k}$, $d_{\pi_{k}}$ is the state distribution induced by the policy $\pi_{k}$, $\delta$ is the target KL divergence, $\gamma$ is the discount factor, $i$ is the cost constraint index number ($i\in\{1\}$ in our experiments, single cost is considered).
@@
To project the optimal update policy $\pi^{*}$ to the parameterized policy $\pi_{\theta}$, minimize the loss function
$\mathcal{L}(\theta)=\underset{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s\sim d_{\pi_{\theta_{k}}}}}{\mathbb{E}}\left[D_{KL}\left(\pi_{\theta}\|\pi^{*}\right)[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}]\right].$ (8)
@@
$\nabla_{\theta}\mathcal{L}(\theta)=\underset{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s\sim d_{\pi_{\theta_{k}}}}}{\mathbb{E}}\left[\nabla_{\theta}D_{KL}\left(\pi_{\theta}\|\pi^{*}\right)[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}]\right]$
$=\nabla_{\theta}D_{KL}\left(\pi_{\theta}\|\pi_{\theta_{k}}\right)[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}]$-
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{\alpha}}\underset{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s\sim d_{\pi_{\theta_{k}}}}}{\mathbb{E}}\left[\sum_{i=0}^{T-1}\frac{\nabla_{\theta}\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}\left(A_{R}^{\pi_{\theta_{k}}}(s_{i},a_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}A_{C}^{\pi_{\theta_{k}}}(s_{i},a_{i})\right)\right].$ (9)
We refer readers to Appendix B of [20] for detailed proof. In implementation we estimate the expectation in Equation 8 and Equation 3.2 by averaging the per-step loss along the most recent on-policy rollout rather than sampling isolated states from a buffer. Partial observability and the history-dependent submodular reward require computing the recurrent hidden state (more in Section 3.4) along a contiguous trajectory, and reconstructing them for random samples would require re-encoding full episodes after each policy update. This full trajectory "sampling" does not violate the validity of policy loss of FOCOPS.
Let $\tau=(s_{0},a_{0},s_{1},a_{1},...,s_{T-1},a_{T-1})$ represent a trajectory of length $T$, the reward advantage $A_{R}^{\pi_{\theta_{k}}}(s_{i},a_{i})$ is defined in a submodularity-inspired trajectory-wise way, named Marginal Gain Advantage Estimation (MGAE) for reward,
@@
where $\Delta(s_{j+1}\mid\tau_{0:j})\equiv\Delta(s_{j+1}\mid S_{j})$ represents the true marginal gain of the state $s_{j+1}$ given the trajectory $\tau_{0:j}$ up to horizon $j$ (Equation 3), and $b(\tau_{0:i})$ is the baseline function dependent on the trajectory up to horizon $i$ that helps to reduce the variance of reward advantage estimation [62]. The important part here is how to choose the baseline function. It should represent a reasonable anticipation of the future cumulative rewards that can guide the policy improvement by comparing with the actual partial return to form the reward advantage. As proved in [7], baseline can be any history-dependent function. We choose this baseline function to be the agent’s current best guess of the remaining rewards along the current trajectory to further reduce the variance,
$b(\tau_{0:i})=\underset{\tau\sim\pi_{\theta}}{\operatorname{mean}}({\sum_{i=0}^{T-1}\Delta(s_{i+1}\mid\tau_{0:i})})-\sum_{i=0}^{j}\hat{\Delta}(s_{i+1}\mid\tau_{0:i}).$ (11)
The left part in Equation 11 is the mean trajectory return across a sliding window of trajectories up to the baseline calculation ($\tau\sim\pi_{\theta}$), whereas the right part is the sum of estimated marginal gains up to step $j$ along the current trajectory. $\hat{\Delta}(s_{i+1}\mid\tau_{0:i})$ is the estimated marginal gain of the state $s_{i+1}$ given the trajectory $\tau_{0:i}$ up to horizon $i$. MGAE uses a history-aware, action-independent baseline aligned with the marginal-gain structure. It is one practical choice among many unbiased baselines. We select it for simplicity and for direct correspondence to the submodular objective. Exploring alternative baselines is promising in future work. Figure 1 contrasts MGAE, which leverages marginal gains, with GAE [9], which relies on a value critic.
@@ 3.3 Semantic Dynamics Model
To further improve spatial consistency, our method employs a geometrically constrained semantic dynamics model (SDM), where a neural network estimates the homography between consecutive semantic observations. This approach assumes that a significant portion of pixels in the image represents a flat surface (e.g., water) in the 3D environment, allowing the homography to approximate the geometric transformation between frames effectively. Additionally, the model assumes that most objects in the scene are static, which helps simplify the estimation process. By preserving the spatial transformations under these assumptions, our method enables the RL model to better capture short-term dynamic changes in the environment.